# Variational Learning Finds Flatter Solutions at the Edge of Stability

### v2

@@ Abstract
Variational Learning (VL) has recently gained popularity for training deep neural networks. Part of its empirical success can be explained by theories such as PAC-Bayes bounds, minimum description length and marginal likelihood, but there are few tools to unravel the implicit regularization in play. Here, we analyze the implicit regularization of VL through the Edge of Stability (EoS) framework. EoS has previously been used to show that gradient descent can find flat solutions and we extend this result to VL to show that it can find even flatter solutions. This is obtained by controlling the posterior covariance and the number of Monte Carlo samples from the posterior. These results are derived in a similar fashion as the standard EoS literature for deep learning, by first deriving a result for a quadratic problem and then extending it to deep neural networks. We empirically validate these findings on a wide variety of large networks, such as ResNet and ViT, to find that the theoretical results closely match the empirical ones. Ours is the first work to analyze the EoS dynamics in VL.
@@ 1 Introduction
Variational Learning (VL) has been used to perform deep learning from early on (Graves, 2011; Blundell et al., 2015) and recently also started to show good results at large scale. It can now outperform state-of-the-art optimizers without any increase in the cost. For example, on ImageNet, VL substantially improves overfitting commonly seen in AdamW and for pretraining GPT-2 from scratch, VL achieves a lower validation perplexity than AdamW (Shen et al., 2024). For low-rank fine-tuning of Llama-2 (7B), VL improves both accuracy (by $2.8\%$) and calibration (by $4.6\%$) (Cong et al., 2024; Li et al., 2025). Moreover, VL methods explicitly derived from minimizing PAC-Bayes bounds (Wang et al., 2023b; Zhang et al., 2024b) have shown consistent improvements over AdamW. All such results confirm the importance of VL for deep learning.
@@
Despite these successes, the theoretical mechanisms behind the good performance of VL remain poorly understood. It is often assumed that simplistic Gaussian posteriors, such as those used currently for deep learning, may not be enough because they are poor approximations of the true posterior; lack of a good prior is another issue. Despite these concerns, VL shows good performance in practice. Theories such as minimum description length (Hinton and Van Camp, 1993; Hochreiter and Schmidhuber, 1997; Blier and Ollivier, 2018), PAC-Bayes (Dziugaite and Roy, 2017; Zhou et al., 2019; Lotfi et al., 2024) and marginal likelihood (Smith and Le, 2017; Immer et al., 2021) can partially explain the success. In particular, PAC-Bayes theory provides a natural explanation for why flatter solutions may generalize better, see for instance the discussion in (Alquier et al., 2024, Section 3.3). However these theories say little about the regularizing properties of the learning algorithm. Similarly to deep learning, the presence of implicit regularization is likely also at play in VL, but few tools exist (if any) to unravel these effects.
@@
In this work, we analyze the implicit regularization of VL algorithms at the EoS. The EoS analysis has previously been used to show that Gradient Descent (GD) with constant learning rate $\rho$ implicitly biases the trajectories towards flatter solutions, where the sharpness (defined as the operator norm of the loss Hessian) hovers around $2/\rho$. Here, we extend this analysis to VL and show that the sharpness can be made lower by controlling the posterior covariance and the number of Monte-Carlo samples used to compute posterior expectations. Similarly to the standard EoS technique, we first derive an exact expression for the stability threshold for VL on a quadratic problem (Figure  1a) and then propose extensions to general loss functions (Figure  1b). We then empirically validate these finding on a wide variety of networks, including Multi-Layer Perceptron, ResNet, and Vision Transformers. We observe similar results when posterior shape is automatically learned, for instance, by using diagonal covariance Gaussians and heavy-tailed posteriors.
@@
Figure 1: (a) The left figure shows trajectory traces of VL on a quadratic problem with an isotropic posterior whose mean is learned but variance is fixed. The trajectory becomes more unstable as the variance is increased (columns) and number of Monte-Carlo samples are decreased (rows). We provide an exact expression to compute the stability threshold at which the iterations become unstable (Theorem 5). We show the validity of the threshold on neural network training. (b) The right figure (top) shows this on CIFAR-10 for an MLP where VL achieves lower sharpness than GD by increasing the posterior variance. The bottom figure shows that the sharpness (solid line) matches the stability threshold obtained by our theorem (dashed line).
@@ 2 Theoretical Tools to Analyze Variational Learning
This objective is equivalent to minimizing the KL divergence to the Gibbs posterior and naturally encourages higher-entropy (flatter) solutions due to the entropy $\mathcal{H}(q)$; for example, see the illustrative example in Khan and Rue (2023, Figure 1). Similar arguments can also be made with the minimum-description-length principles (Hinton and Van Camp, 1993; Graves, 2011; Blier and Ollivier, 2018). In practice, Variational Learning has started to show good results achieving generalization performance better than the state of the art optimizers across several tasks (Shen et al., 2024; Cong et al., 2024). This matches with the intuition: because the variational objective prefers wider distributions, we expect the posterior to be located in the region where the loss $\ell(\boldsymbol{\theta})$ is flatter. Despite this intuition, there is little work to analyze the implicit regularization that could help us understand how VL favors flatter regions and how we can control it. The answer could lie in the shape of the posterior but currently there are no results explicitly characterization the properties of the posterior responsible for improved generalization.
Variational learning is also closely connected to weight-noise or weight-perturbation methods where the goal is to optimize $\mathbb{E}_{q(\boldsymbol{\epsilon)}}[\ell(\boldsymbol{\theta}+\boldsymbol{\epsilon)}]$ where $q(\boldsymbol{\epsilon)}$ is a fixed distribution to inject weight noise. This can be seen as a special case of variational learning where the shape of the posterior is fixed and only the location is learned. Multiple works (Zhu et al., 2019; Nguyen et al., 2019; Zhang et al., 2019; Jin et al., 2017; Simsekli et al., 2019) have analyzed generalization behavior of such weight-noise/variational methods but, even for this simple case, there are no studies connecting them to the EoS results for GD. In this paper, we will address these gaps and provide both theoretical and empirical results regarding the edge of stability phenomenon in variational GD.
@@ 2.1 Edge of Stability for Gradient Descent
To understand how weight-perturbed Gradient Descent (GD) operates at the Edge of Stability (EoS), we first explain the main result regarding the behavior of GD on a quadratic loss. Afterward, we discuss its implication for deep neural networks. We consider the quadratic loss shown below:
[…]
@@
(Descent Lemma) Consider the GD update $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\rho\nabla\ell(\boldsymbol{\theta}_{t})$ on the quadratic loss (2). The loss decreases at each step as
@@
This lemma implies that GD converges on a quadratic loss if all eigenvalues satisfy $\lambda_{i}<2/\rho$. This is a different way of writing the standard condition that maximum eigenvalue $\lambda_{1}<2/\rho$. The result extends to any smooth $\ell(\boldsymbol{\theta})$ and for such case, the condition implies that the learning rate $\rho<2/\beta$, where $\beta=\sup_{\boldsymbol{\theta}}\|\nabla^{2}\ell(\boldsymbol{\theta})\|_{2}$ is the bound on the Hessian norm (Nesterov et al., 2018, Chapter 2). This condition is necessary and sufficient for descent of GD on quadratic.
@@
While the descent lemma is predictive of convergence of GD for smooth functions, deep learning exhibits a more complex behavior. When training deep neural networks with a constant learning rate $\rho$, the Hessian’s operator norm (or sharpness) tends to settle around the value $2/\rho$. This phenomenon is referred to as the Edge of Stability. Deep neural networks often operate at this boundary (hence the nomenclature) and converge in a non-monotonic, unstable manner. A comprehensive empirical investigation of this phenomenon at least dates back to Cohen et al. (2021), where two phases were noticed when training neural networks. In the first phase referred to as ‘progressive sharpening’, the Hessian’s operator norm, $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$ increases and slowly approaches $2/\rho$. This is followed by the second, the EoS phase, where sharpness hovers around $2/\rho$ and the loss continues to decrease in an oscillatory fashion. Increasing $\rho$ reduces the sharpness $2/\rho$, which could then drive the iterates toward flatter minima that may generalize better. EoS analysis can help to understand such implicit regularization during training.
@@
EoS is especially useful for nonconvex problems, such as those in deep learning, as it helps narrow down the region of the loss landscape where the iterates tend to converge. In its early phase, neural network training behaves like optimization on a quadratic loss, the loss decreases monotonically, and the sharpness $\|\nabla^{2}\ell(\boldsymbol{\theta})\|_{2}$ gradually increases to $2/\rho$. Once the sharpness reaches $2/\rho$, a quadratic approximation would predict that the loss should diverge. However, as shown by Damian et al. (2023), once sharpness reaches $2/\rho$, a local quadratic approximation is insufficient to capture the dynamics because the third order Taylor expansion term of the loss becomes significant. This cubic term represents the gradient of the sharpness, which serves as a negative feedback to counteract progressive sharpening and stabilizes the sharpness around $2/\rho$ in GD. This phenomenon does not happen on a quadratic loss, but only for certain non-convex losses with nonzero third-order derivative. Then the iterates exhibit oscillatory or non-monotonic behavior (instead of divergence) even when $\|\nabla^{2}\ell(\boldsymbol{\theta}_{*})\|_{2}\approx 2/\rho$. This is the reason why $2/\rho$ is also called the ‘Stability Threshold’.
@@
Different optimizers have different stability threshold which depends on the sharpness value $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$. For instance, Sharpness Aware Minimization (SAM) leads to a different, smaller Stability Threshold (Long and Bartlett, 2024) compared to $2/\rho$ for GD. In this work, we show a similar result where VL has a smaller threshold than GD. We show this by deriving the Stability Threshold for a simpler quadratic problem first and analyze several factors such as posterior covariance and the number of posterior samples that influence the threshold. Then, we empirically demonstrate that similar results hold for the case when VL is used to train deep neural networks.
@@ 3 Stability Threshold for A Simple Case of Variational Learning
We start with a simple VL setting where the goal is to estimate a Gaussian $q(\boldsymbol{\theta})=\mathcal{N}(\boldsymbol{\theta}|\,\boldsymbol{m},\boldsymbol{\Sigma})$ whose mean $\boldsymbol{m}$ is unknown and covariance $\boldsymbol{\Sigma}$ is fixed. We assume $\boldsymbol{\Sigma}$ is a diagonal covariance matrix, with $\sigma^{2}_{i}$ representing its $i^{th}$ largest value. Khan and Rue (2023) [Section 1.3.1] show that this can be estimated by the following Variational GD algorithm:
@@
$\mathbf{m}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})}[\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon})]$ (4)
@@
This algorithm can be implemented by the following version where the expectation is estimated using $N_{s}$ Monte Carlo samples as follows
@@
$\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon_{i}}).\quad\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$ (5)
@@
The updated iterate $\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}$, conditioned on $\mathbf{m}_{t}$, depends on the $N_{s}$ Monte Carlo samples $\boldsymbol{\epsilon}=[\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2},..,\boldsymbol{\epsilon}_{N_{s}}]$. Compared to standard gradient descent, the update step in Variational GD, is determined by gradients averaged over a local neighborhood of perturbed weights. As a result, the update introduces two interacting effects influencing its stability threshold: (1) a perturbation effect, originating from the perturbation covariance $\boldsymbol{\Sigma}$, and (2) a smoothing effect, resulting from averaging gradients across $N_{s}$ Monte Carlo samples. Similar to Lemma 2.1, we can derive the stability threshold for the above Variational GD and show that it is smaller than that of GD. {restatable}theoremmainproof Consider the Variational GD update (5) on the quadratic loss (2), then the loss decreases on average as
@@
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]-\ell(\mathbf{m}_{t})<0\quad\text{if}\quad\lambda_{i}<\frac{2}{\rho}\cdot\operatorname{VF}\left(\frac{N_{s}}{\sigma^{2}_{i}}\cdot c_{i,t}\right)\quad\text{for all }i.$ (6)
where $c_{i,t}=(\lambda_{i}\mathbf{m}_{t}^{\top}\mathbf{v}_{i})^{2}$, and $\operatorname{VF}(\cdot)$ denotes the Variational Factor (7), a monotonically increasing function satisfying $\operatorname{VF}(\cdot)<1$. See Appendix B for a proof. The theorem is analogous to the descent lemma for GD and it states that the Variational GD (5) also decreases the expected loss over $\boldsymbol{\epsilon}$ if each eigenvalue $\lambda_{i}$ is less than $2/\rho$ times a function called the Variational Factor (VF). The VF function is strictly less than $1$, therefore the stability threshold is strictly less than $2/\rho$. Unlike GD, the condition here is only sufficient but not necessary. A necessary condition can also be derived but we find the one above to be sufficient for our empirical results. Further details are along with extensive experiments are given in Appendix E.
An exact expression for the VF factor is given as:
@@
$\text{VF}(z):=\rho\cdot\sqrt{\frac{z}{3}}\cdot\sinh\left(\frac{1}{3}\operatorname{arcsinh}\left(\frac{3}{\rho}\sqrt{\frac{3}{z}}\right)\right).$ (7)
@@
for a scalar $z>0$. Figure 2(a) shows the stability threshold as a function of $z$ where it is clear that it always lower bounds $2/\rho$. The exact value of VF depends on its argument $z$ which in eq. 5 mainly depends on the number of Monte-Carlo samples $N_{s}$ and posterior variance $\sigma^{2}_{i}$, but also on the loss-dependent constant $c_{it}$ which is essentially a function of $\lambda_{i}$ and $\mathbf{m}_{t}^{\top}\mathbf{v}_{i}$.
@@
Theorem 1 guarantees a decrease in the expected loss. Below, we state another result to show that the actual loss also decreases with high probability if the expected loss decreases by a margin $\delta>0$. {restatable}lemmadescentconc In the same setting as 5, when the loss on average, decreases by some margin $\delta>0$:
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]-\ell(\mathbf{m}_{t})<-\delta\quad\text{then,}$,
$\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})-\ell(\mathbf{m}_{t})<0\;\text{occurs with probability at least}\;1-2\exp\left(-c_{1}\min\left\{\delta^{2}N_{s}^{2}/c_{2},\delta N_{s}/c_{2}\right\}\right),$
@@
for constants $c_{1},c_{2}>0$ depending only on $\rho$, $\mathbf{Q}$, and $\boldsymbol{\Sigma}$. The result also shows that the probability with which this happens also increases with the number of Monte Carlo samples $N_{s}$.

### v3

@@ Abstract
Variational Learning (VL) has recently gained popularity for training deep neural networks. Part of its empirical success can be explained by theories such as PAC-Bayes bounds, minimum description length and marginal likelihood, but little has been done to unravel the implicit regularization in play. Here, we analyze the implicit regularization of VL through the Edge of Stability (EoS) framework. EoS has previously been used to show that gradient descent can find flat solutions and we extend this result to show that VL can find even flatter solutions. This result is obtained by controlling the shape of the variational posterior as well as the number of posterior samples used during training. The derivation follows in a similar fashion as in the standard EoS literature for deep learning, by first deriving a result for a quadratic problem and then extending it to deep neural networks. We empirically validate these findings on a wide variety of large networks, such as ResNet and ViT, to find that the theoretical results closely match the empirical ones. Ours is the first work to analyze the EoS dynamics of VL.
@@ 1 Introduction
Variational Learning (VL) has been used to perform deep learning from early on (Graves, 2011; Blundell et al., 2015) and recently also started to show good results at large scale. It has been shown to outperform state-of-the-art optimizers without any increase in the cost. For example, on ImageNet, VL substantially improves overfitting commonly seen in AdamW and for pretraining GPT-2 from scratch, VL achieves a lower validation perplexity than AdamW (Shen et al., 2024). For low-rank fine-tuning of Llama-2 (7B), VL improves both accuracy (by $2.8\%$) and calibration (by $4.6\%$) (Cong et al., 2024; Li et al., 2025). Moreover, VL methods explicitly derived by using PAC-Bayes bounds (Wang et al., 2023b; Zhang et al., 2024b) have shown consistent improvements over AdamW. All such results confirm the importance of VL for deep learning.
@@
Despite these successes, the theoretical mechanisms behind the good performance of VL remain poorly understood. It is often assumed that simplistic Gaussian posteriors, such as those used currently for deep learning, may not be enough because they are poor approximations of the true posterior; lack of a good prior is another issue. Despite these concerns, VL shows good performance in practice. Theories such as minimum description length (Hinton and Van Camp, 1993; Hochreiter and Schmidhuber, 1997; Blier and Ollivier, 2018), PAC-Bayes (Dziugaite and Roy, 2017; Zhou et al., 2019; Lotfi et al., 2024; Alquier et al., 2024) and marginal likelihood (Smith and Le, 2017; Immer et al., 2021) can partially explain the success. In particular, PAC-Bayes theory provides a natural explanation for why flatter solutions may generalize better, see for instance the discussion in (Alquier et al., 2024, Section 3.3). However these theories say little about the regularizing properties of the learning algorithm. Similarly to deep learning, the presence of implicit regularization is likely also at play in VL, but few tools exist to unravel these effects.
@@
In this work, we analyze the implicit regularization of VL algorithms at the Edge of Stability. The EoS analysis has previously been used to show that Gradient Descent (GD) with constant learning rate $\rho$ implicitly biases the trajectories towards flatter solutions, where the sharpness (defined as the operator norm of the loss Hessian) hovers around $2/\rho$. We extend this analysis to VL and show that sharpness can be further lowered by controlling the posterior covariance and the number of Monte-Carlo samples used to compute posterior expectations; see an illustration in Figure 1(a). Similarly to the standard EoS technique, we first derive an exact expression for the stability threshold for VL on a quadratic problem and then propose extensions to general loss functions. We then empirically validate these finding on a wide variety of deep networks, including Multi-Layer Perceptron, ResNet, and Vision Transformers; see an example in Figure 1(b). We observe similar results when posterior shape is automatically learned, for instance, by using diagonal covariance Gaussians and heavy-tailed posteriors. Code to replicate these results is available at https://github.com/Avra98/variationallearning-eos.
@@
Figure 1: Panel (a): The left figure shows trajectory traces of VL on a quadratic problem with an isotropic variational posterior whose mean is learned but variance is set to a fixed value. The trajectory becomes more unstable as the posterior variance is increased and number of Monte-Carlo samples is decreased. We provide an exact expression to compute the stability threshold at which the iterations become unstable (Theorem 5). Panel (b): We show the validity of the threshold on neural network training. The right figure (top) shows this on CIFAR-10 for an MLP where VL achieves lower sharpness than GD when posterior variance is increased. The bottom figure shows that the sharpness (solid line) matches the stability threshold obtained by our theorem (dashed line).
@@ 2 Theoretical Tools to Analyze Variational Learning
This is an instance of the maximum-entropy principle. The objective is equivalent to minimizing the KL divergence to the Gibbs posterior and naturally encourages higher-entropy (flatter) solutions due to the entropy $\mathcal{H}(q)$; for example, see the illustrative example in Khan and Rue (2023, Figure 1). Similar arguments can also be made with the minimum-description-length principles (Hinton and Van Camp, 1993; Graves, 2011; Blier and Ollivier, 2018). In practice, Variational Learning has started to show good results achieving generalization performance better than the state of the art optimizers across several tasks (Shen et al., 2024; Cong et al., 2024). This matches with the intuition: Because the variational objective prefers wider distributions, we expect the posterior to be located in the region where the loss $\ell(\boldsymbol{\theta})$ is flatter, see Figure 2.
Despite this intuition, there is little work to analyze the implicit regularization that could help us understand how VL favors flatter regions and how we can control it. Existing theories do not sufficiently address this; a review of the related work is in Appendix A. We suspect the implicit regularization to be related to the shape of the posterior but currently there are no results explicitly characterizing this.
Figure 2: VL’s mechanism for flatter minima: The posterior variance determines the minima’s location. A small variance settles the posterior in a sharp minima (left), while a larger variance allows it to explore and find a flat minima (right).
VL is also closely connected to weight-noise or weight-perturbation methods where the goal is to optimize $\mathbb{E}_{q(\boldsymbol{\epsilon)}}[\ell(\boldsymbol{\theta}+\boldsymbol{\epsilon)}]$ with $q(\boldsymbol{\epsilon)}$ being a fixed distribution to inject weight noise. This can be seen as a special case of variational learning where the shape of the posterior is fixed and only the location is learned. Multiple works (Zhu et al., 2019; Nguyen et al., 2019; Zhang et al., 2019; Jin et al., 2017; Simsekli et al., 2019) have analyzed generalization behavior of such weight-noise variational methods but, even for this simple case, there are no studies connecting them to the EoS results for GD. In this paper, we will address these gaps and provide both theoretical and empirical results regarding the edge of stability phenomenon of such a variational GD (VGD).
@@ 2.1 Edge of Stability for Gradient Descent
We will briefly review the EoS result for GD. The standard EoS literature relies on a result for a quadratic problem and then extends it to deep neural networks. For instance, consider the following
$\text{quadratic loss:}\quad\ell(\boldsymbol{\theta})=\frac{1}{2}\boldsymbol{\theta}^{\top}\mathbf{Q}\boldsymbol{\theta},\quad\text{where}\quad\mathbf{Q}=\sum_{i=1}^{d}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{T}_{i}.$ (2)
@@
(Descent Lemma) For a GD update $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\rho\nabla\ell(\boldsymbol{\theta}_{t})$ on the quadratic loss (2), the loss decreases at each step, that is, we have
@@
This lemma implies that GD converges on a quadratic loss if all eigenvalues satisfy $\lambda_{i}<2/\rho$. This is a different way of writing the standard condition that maximum eigenvalue $\lambda_{1}<2/\rho$. The result extends to any smooth $\ell(\boldsymbol{\theta})$ and for such cases, the condition implies that the learning rate $\rho<2/\beta$, where $\beta=\sup_{\boldsymbol{\theta}}\|\nabla^{2}\ell(\boldsymbol{\theta})\|_{2}$ is the bound on the Hessian norm (Nesterov et al., 2018, Chapter 2). This condition is necessary and sufficient for descent of GD on quadratic.
@@
While the descent lemma is predictive of convergence of GD for smooth functions, deep learning exhibits a more complex behavior. When training deep neural networks with a constant learning rate $\rho$, the Hessian’s operator norm (or sharpness) tends to settle around the value $2/\rho$. This phenomenon is referred to as the Edge of Stability. Deep neural networks often operate at this edge and converge in a non-monotonic, unstable manner. A comprehensive empirical investigation of this phenomenon is given by Cohen et al. (2021), where two phases of training neural networks were noticed. In the first phase referred to as ‘progressive sharpening’, the Hessian’s operator norm, $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$ increases and slowly approaches $2/\rho$. This is followed by the second, EoS phase, where sharpness hovers around $2/\rho$ and the loss continues to decrease in an oscillatory fashion.
@@
This phenomenon does not happen on a quadratic loss, but only for certain losses with nonzero third-order derivative. As shown by Damian et al. (2023), once sharpness reaches $2/\rho$, a local quadratic approximation is insufficient to capture the dynamics because the third order Taylor expansion term of the loss becomes significant. This cubic term represents the gradient of the sharpness, which serves as a negative feedback to counteract progressive sharpening and stabilize the sharpness around $2/\rho$ in GD. Instead of divergence, the iterates exhibit oscillatory or non-monotonic behavior even when the sharpness reaches $2/\rho$. This is the reason why $2/\rho$ is also called the ‘Stability Threshold’. Increasing $\rho$ reduces the edge, which could then drive the iterates toward flatter minima that may generalize better. EoS analysis can help us understand such implicit regularization during training, especially for nonconvex problems.
@@
Different optimizers have different stability thresholds which depends on the sharpness value $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$. For instance, Sharpness Aware Minimization (SAM) leads to a different, smaller stability threshold (Long and Bartlett, 2024) compared to $2/\rho$ for GD. In this work, we show a similar result where VL has a smaller threshold than GD. We show this by deriving the stability threshold for a simpler quadratic problem first and analyze several factors such as posterior covariance and the number of posterior samples that influence the threshold. Then, we empirically demonstrate that similar results hold for the case when VL is used to train deep neural networks.
@@ 3 Stability Threshold for a Simple Case of Variational Learning
We start with a simple VL setting where the goal is to estimate a Gaussian $q(\boldsymbol{\theta})=\mathcal{N}(\boldsymbol{\theta}|\,\boldsymbol{m},\boldsymbol{\Sigma})$ whose mean $\boldsymbol{m}$ is unknown and covariance $\boldsymbol{\Sigma}$ is fixed. We assume $\boldsymbol{\Sigma}=\sigma^{2}\mbox{$\mbox{$\mathbf{I}$}$}$ is an isotropic covariance matrix, with scalar variance $\sigma^{2}$. Khan and Rue (2023) [Section 1.3.1] show that this can be estimated by the following Variational GD algorithm:
@@
$\mathbf{m}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})}[\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon})].$ (4)
@@
This algorithm can be implemented by the following version where the expectation is estimated by drawing $N_{s}$ Monte Carlo samples to approximate the expectation as follows
@@
$\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon_{i}}),\quad\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mbox{$\mbox{$\mathbf{I}$}$}).$ (5)
@@
The updated iterate $\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}$ depends on the $N_{s}$ Monte Carlo samples $\boldsymbol{\epsilon}=[\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2},..,\boldsymbol{\epsilon}_{N_{s}}]$. Compared to the standard gradient descent, the update step in Variational GD (VGD), is determined by gradients averaged over a local neighborhood of perturbed weights. As a result, the update introduces two interacting effects influencing its stability threshold: (1) a perturbation effect, originating from the perturbation covariance $\boldsymbol{\Sigma}$, and (2) a smoothing effect, resulting from averaging gradients across $N_{s}$ Monte Carlo samples. Similarly to Lemma 2.1, we can derive the stability threshold for the above VGD and show that it is smaller than that of GD. {restatable}theoremmainproof Consider the VGD update (5) on the quadratic loss (2), then
@@
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]-\ell(\mathbf{m}_{t})<0\quad\text{if}\quad\lambda_{i}<\frac{2}{\rho}\cdot\operatorname{VF}\left(\frac{N_{s}}{\sigma^{2}}\cdot c_{i,t}\right)\quad\text{for all }i,$ (6)
where $c_{i,t}=(\lambda_{i}\mathbf{m}_{t}^{\top}\mathbf{v}_{i})^{2}$, and $\operatorname{VF}(\cdot)$ denotes the Variational Factor given by
@@
$\operatorname{VF}(z):=\rho\cdot\sqrt{\frac{z}{3}}\cdot\sinh\left(\frac{1}{3}\operatorname{arcsinh}\left(\frac{3}{\rho}\sqrt{\frac{3}{z}}\right)\right).$ (7)
See Appendix B for a proof where we also discuss the case where $\mathbf{Q}$ is low rank. The theorem is analogous to the descent lemma for GD and it states that the Variational GD (5) also decreases the expected loss over $\boldsymbol{\epsilon}$ if each eigenvalue $\lambda_{i}$ is less than $2/\rho$ times a function called the Variational Factor (VF). The VF function is strictly less than $1$, therefore the stability threshold is strictly less than $2/\rho$. Unlike GD, the condition here is only sufficient but not necessary. A necessary condition can also be derived but the one above is sufficient for this paper.
@@
Figure 3(a) shows the stability threshold as a function of $z$ where it is clear that it always lower bounds $2/\rho$ (dashed horizontal line). The exact value of VF depends on its argument $z$ which in Equation˜5 mainly depends on the number of Monte-Carlo samples $N_{s}$ and posterior variance $\sigma^{2}$, but also on the loss-dependent constant $c_{it}$ which is essentially a function of $\lambda_{i}$ and $\mathbf{m}_{t}^{\top}\mathbf{v}_{i}$.
@@
Theorem 1 guarantees a decrease in the expected loss. Below, we state another result to show that the actual loss also decreases with high probability if the expected loss decreases by a margin $\delta>0$. {restatable}lemmadescentconc In the same setting as Equation˜5, when the expected loss at next iteration is smaller than the previous loss by some margin $\delta>0$, that is,
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]<\ell(\mathbf{m}_{t})-\delta,$
@@
then $\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})-\ell(\mathbf{m}_{t})<0$ occurs with probability at least $1-2\exp\left(-c_{1}\min\left\{\delta^{2}N_{s}^{2}/c_{2},\delta N_{s}/c_{2}\right\}\right)$, for constants $c_{1},c_{2}>0$ depending only on $\rho$, $\mathbf{Q}$, and $\boldsymbol{\Sigma}$. The result also shows that the probability with which this happens increases with the number of Monte Carlo samples $N_{s}$.